OpenStudy (anonymous):

sin theta=-2/9 and tan theta is less than 0. Find sec theta and draw a reference angle. PLEASE HELP IM DYING I don't understand this

Nnesha (nnesha):

 remember this CAST rule for signs

OpenStudy (prizzyjade):

Sin theta = -2/9 Tan theta<0 Sec theta = 1/cos theta Tan theta = sintheta/costheta Let tan = -1 -1(costheta)= -2/9 Costheta=2/9 Sectheta=1/(2/9) Sectheta= 9/2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I'm still so confused haha

Nnesha (nnesha):

alright so given sin = -2/9 which is negative 2nd information given is tan < 0 so at which quadrant you should draw right triangle ?